OpenStudy (mphetang):

what are the formal charges of the elements in the phosphite ion, PO3-3

OpenStudy (anonymous):

F.C (P) = 5 - (1/2 * 6 + 2) = 0 F.C (O) = 6 - (1/2 * 2 + 6) = -1